OpenStudy (anonymous):

Solve this equation: sin(x + (π / 4)) – sin(x – (π / 4) = 1

OpenStudy (sidsiddhartha):

use sin(a+b)=sina*cosb+cosa*sinb sin(a-b)=sina*cosb-cosa*sinb

OpenStudy (anonymous):

can you show me?

OpenStudy (sidsiddhartha):

sin(x+pi/4)=sinx*cos(pi/4)+cosx*sin(pi/4) =sinx/sqrt(2)+cosx/sqrt(2)...............1 sin(x-pi/4)=sinx*cos(pi/4)-cosx*sin(pi/4) =sinx/sqrt(2)-cosx/sqrt(2).................2 substract (1)-(2) 2cosx/sqrt(2)=1 cosx=1/sqrt(2) x=pi/4 that's ur answer

OpenStudy (anonymous):

oh ok I follow that. is that the only answer or would there be another one?

OpenStudy (sidsiddhartha):

i think thats the one

OpenStudy (anonymous):

@Hero just to verify. @sidsiddhartha explained this really well

OpenStudy (anonymous):

how did you go from 2cosx / sqrt2 = 1 to cosx=1/sqrt2. what happened to the 2 in 2cosx

OpenStudy (anonymous):

@sidsiddhartha

OpenStudy (sidsiddhartha):

2/sqrt(2)=sqrt(2)

OpenStudy (anonymous):

thanks man! hopefully that is the only answer

OpenStudy (anonymous):

@Hero can you just recheck this? wolfram doesn't agree but i may have entered it wrong

hero (hero):

@washcaps, at some point, you have to be able to think yourself in this regard. He showed you the steps. If you understood it, and checked the answer, then there either is a mistake or there isn't one. If there is a mistake, you should be competent enough (if you understand the material) to go and locate where the mistake is. If there is no mistake, then you should check how you're inputting the information into wolfram. One way to check to see if a mistake was made is by graphing the trigonometric expression. For that, I recommend http://www.desmos.com

hero (hero):

If we do all of the work for you, then you won't have learned anything useful for the next set of problems you'll be tasked with.

OpenStudy (anonymous):

RIght. The equation and work looked fine to me, but wolfram didn't agree, so I don't know what to do. I can't find an error based off of what I know

hero (hero):

I'm just saying, if it were me, I would be asking myself Did I make a mistake while verifying the result? Is there an algebraic mistake in the written solution? Does the graph of the expression accurately reflect the given solution?

OpenStudy (anonymous):

based off of my knowledge there was no mistake, but you seem to know this material, so I was just asking you if there was a problem

hero (hero):

I showed you as much as I'm willing to show you for today

OpenStudy (anonymous):

can you not just tell me if the answer is right or wrong? If it's right (which I think it is) then great. If its wrong, then I need to figure out why

OpenStudy (anonymous):

@Hero can you help now?

OpenStudy (anonymous):

@RadEn

hero (hero):

You've been waiting here for three hours?

hero (hero):

I figure that maybe you would have figured it out by now. All you have to do is apply the rule.

hero (hero):

Have you tried doing it on your own. Applying the formula takes only a little bit of effort. As I said before, no one wants to do your work for you.